Use the principle of separation of variables to solve dy xyx0 dx Find the constant of integration for y(0) = 1.
Answer #1

Solution: ( = 0 4 - da vy += 0 se on a rey-x → ter 2(4-1) seorang ada Integrate both sides Som fx dx = lng-) = 42+c = yta ezt